OpenStudy (geekfromthefutur):

Two sets of quadratic expressions are shown below in various forms: Line 1: x2 − x − 6 (x + 2)(x − 3) (x + 0.5)2 − 6.25 Line 2: x2 − x − 2 (x + 1)(x − 2) (x − 0.5)2 − 2.25 Which line contains three equivalent expressions? Line 1 only Line 2 only Both Line 1 and Line 2 Neither Line 1 nor Line 2

OpenStudy (geekfromthefutur):

@Nurali

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

do this simplify x^2-x-6 and x^2-x-2

OpenStudy (anonymous):

in () form

OpenStudy (anonymous):

you will get (x+2)(x-3) for first and for 2nd (x + 1)(x − 2)

OpenStudy (anonymous):

so atleast till this point they are equivalent

OpenStudy (geekfromthefutur):

yea :D

OpenStudy (anonymous):

well now we have to see for 3rd

OpenStudy (anonymous):

now put some value for x suppose x=2

OpenStudy (anonymous):

when you solve (x+2)(x-3) =-4 and (x + 1)(x − 2) = 0

OpenStudy (anonymous):

hey is that squared it last equation

OpenStudy (geekfromthefutur):

i think so im not quit sure

OpenStudy (anonymous):

so for x=2 we get (x + 0.5)2 − 6.25 = 0 and (x − 0.5)2 − 2.25 = 4

OpenStudy (anonymous):

the values are not equal

OpenStudy (anonymous):

so neither of them have equivalent expression

OpenStudy (geekfromthefutur):

ok thank you so much :3

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

did u get this right?